OpenStudy (anonymous):

Find the domain G(x) = x+4/x^3-4x Please show all work.

OpenStudy (anonymous):

well I'm not sure if there's more to it, but obviously denominator can't equal 0. So, x^3-4x=0 x(x^2-4)=0 x(x+2)(x-2)=0 x=-2,0,2 I think other than these three numbers, there are no restrictions on domain.

OpenStudy (anonymous):

Thank you